In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?
Answer by josmiceli(6797)

w = winners votes
l = losers votes
w = l + 220
l + w = 810
l + l + 220 = 810
2l = 810 - 220
2l = 590
l = 295 losers votes
w = l + 220
w = 295 + 220
w = 515 winners votes
check
l + w = 810
295 + 515 = 810
810 = 810
OK
l + l + 220 = 810
590 + 220 = 810
810 = 810
OK